OpenStudy (anonymous):

2x^2 - x + 10 = 0 HELP ME PLEASE!!! I will give a medal!!!!

OpenStudy (solomonzelman):

use the quadratic formula.

OpenStudy (anonymous):

yeah, I've got it now thanks!!!

OpenStudy (anonymous):

I gave you a medal just for helping me!!! :)

OpenStudy (solomonzelman):

yeah, that was really kind :) Do it and show me what you get, to make sure you're correct.

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

hold on, are you familiar with \(\color{blue}{ i }\) ?

OpenStudy (anonymous):

????

OpenStudy (solomonzelman):

No? Because you will get a negative root if you do the quadratic formula.

OpenStudy (anonymous):

here are the solutions!! -2, 2.5 -1.97, 2.47 -2.5, 2 no solution

OpenStudy (solomonzelman):

\(\color{blue}{ i =\sqrt{-1}}\)

OpenStudy (solomonzelman):

Which course are you taking ?

OpenStudy (anonymous):

Algebra 1

OpenStudy (solomonzelman):

If you are taking trig, then say \(\huge\color{blue}{ \frac{2±i\sqrt{79}}{4} }\) but for algebra 1, it would be just no solution, because negative root is no solution (for alg 1 )

OpenStudy (solomonzelman):

yeah. it would be no solush.

OpenStudy (anonymous):

oh ok thank you!!!

OpenStudy (solomonzelman):

You welcome !

OpenStudy (anonymous):

can you help me with this one?

OpenStudy (solomonzelman):

Sure

OpenStudy (anonymous):

-5 + 2x^2 = -6x (what expression gives the solution?)

OpenStudy (solomonzelman):

well, first add 6x to both sides then order in a polynomial degree x^2 then x then the number (the constant)

OpenStudy (anonymous):

\[\frac{ 2\pm \sqrt{4-(4)(6)(-5)} }{ 12 }\]

OpenStudy (anonymous):

^^^^ that is A

OpenStudy (solomonzelman):

you would get \(\color{goldenrod}{ 2x^2+6x-5=0 }\) ooh, working fast on that one !

OpenStudy (solomonzelman):

but look at my equation and then do the quadratic, b/c I think you go things wrong.

OpenStudy (solomonzelman):

2 is not B. you got confused, b/c you didn't put the equation in order. See what I mean?

OpenStudy (anonymous):

\[\frac{ -5\pm \sqrt{25-(4)(2)(6)} }{ -10 }\]

OpenStudy (anonymous):

^^^^^^ that is B, I am having to type them one at a time b/c my computer is slow and lagging!!!

OpenStudy (solomonzelman):

no, incorrect again. HOLD ON! first put the equation in a degree order. you will see what I mean. \(\color{blue}{ -5+2x^2=-6x }\) \(\color{red}{ ~~~+6x~~~~~~+6x }\) \(\color{blue}{ -5+2x^2+6x=0 }\) put this in a normal order \(\color{blue}{ 2x^2+6x-5=0 }\) \(\color{blue}{ a=2~~~~~~b=6~~~~~~c=-5 }\) do it.

OpenStudy (anonymous):

C. \[\frac{ -6\pm \sqrt{36-(4)(2)(-5)} }{ 4 }\] D. \[\frac{ 6\pm \sqrt{36-(4)(2)(5)} }{ 4 }\]

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

knowing your a b c looking a what I said just now (the blue and red stuff in a prev. reply ) which one do you think it is ?

OpenStudy (anonymous):

so it is C?

OpenStudy (solomonzelman):

Yup !

OpenStudy (anonymous):

thanks!!! :D

OpenStudy (anonymous):

I would give you a million medals if I could!!! lol!!! :D

OpenStudy (anonymous):

So for the first question I had, was that quadratic formula?

OpenStudy (solomonzelman):

\(\color{blue}{ Nice~~~~~to~~~~~hear~~~~~that~,~~~~~you~~~~~welcome~! }\)

OpenStudy (solomonzelman):

For the first? well the first question either way was no solution. If you got no solution doing the quadratic formula, you would get no solution using any other method as well.

OpenStudy (anonymous):

would this be quadratic formula 8x^2 - 13x + 3 = 0?

OpenStudy (anonymous):

and ok!!

OpenStudy (solomonzelman):

would you solve this one using the quad. form. ?

OpenStudy (solomonzelman):

that;s your question?

OpenStudy (anonymous):

yes! I guess so!!

OpenStudy (anonymous):

b/c it wouldnt be factoring graphing or square roots!!! lol

OpenStudy (solomonzelman):

yeah, it would be very difficult to complete the square. You CAN'T factor it (into integers, but you could into decimals, but you won't try to do that, b/c ik you aren't stupid) yeah, you would prefer quad. form. as your best method (in this case) .

OpenStudy (anonymous):

Thanks!!! I have just one more question!! :) Could you help me with it?

OpenStudy (solomonzelman):

Yeah, shoot !

OpenStudy (anonymous):

The perimeter of a rectangle is 54 cm. The area of the same rectangle is 176 cm^2. What are the dimensions of the rectangle?

OpenStudy (solomonzelman):

ok, let width be "w" and length be "l" \[P=2(w+l)\] \[A=~w \times l\] substitute what you know for P (Perimeter) and A (Area) and solve. \(\color{blue}{ 54=2(w+l) } \) \(\color{blue}{ 176=w \times l } \)

OpenStudy (solomonzelman):

doesn't matter which one is l and which one is w. (That you won't find) can you solve the system of equation I posted in blue ?

OpenStudy (anonymous):

I can try!!

OpenStudy (anonymous):

yeah, no I got an answer that was way off from my choices!!

OpenStudy (solomonzelman):

Substitute from the first equation into the second. \(\color{blue}{ 54 =2(w+l) } \) \(\color{black}{~~~->~~~} \) \(\color{blue}{ 27 =w+l } \) \(\color{black}{~~~->~~~} \) \(\color{blue}{ 27-w =l } \) substitute \(\color{blue}{ 27-w } \) for \(\color{blue}{ l } \). \(\color{blue}{ 176= w \times l } \) \(\color{blue}{ 176= w \times (27-w) } \) \(\color{blue}{ 176= 27w-w^2 } \) can you solve \(\color{blue}{ 176= 27w-w^2 } \) ?

OpenStudy (anonymous):

176(27) - 176^2?

OpenStudy (solomonzelman):

11 and 16 that's what I got. for w

OpenStudy (anonymous):

oh ok, thank you!!! I really need help in this part of Algebra, good thing my mom is coming home soon!!! Thanks!!!

OpenStudy (anonymous):

Have a great day!!! :) bye!!

OpenStudy (solomonzelman):

yeah, ... \(\color{blue}{ 11\times16=176 ~~~~~~~~~~~:D } \)

OpenStudy (solomonzelman):

yeah, bye !